OpenStudy (anonymous):

1.The problem you created. Remember, the problem must be similar to the inverse variation problems you’ve already studied. You don’t need to show the solution to your problem, but make sure you do know how to solve it so you can tell other students if they have solved it correctly. 2.Your answer to another student’s problem. Make sure you tell in your post what the original problem was that you are answering. >>this is what i answered Y=6 when x=3.find y when x=9

OpenStudy (anonymous):

ok so what do you want me to do???

OpenStudy (anonymous):

I need a simaler problem to y=6 when x=3 . find y when x=9

OpenStudy (anonymous):

you can make y=5 and x=20 but if y=8 then x=??? so x-y=20-5=15 this means x is 15 + y so, x=y+15 x= 15+8 x= 23

OpenStudy (anonymous):

is it ok???

OpenStudy (anonymous):

yes but i just need the equation part

OpenStudy (anonymous):

x= 15+y <=equation part

OpenStudy (anonymous):

So i can put it find y when x=15

OpenStudy (anonymous):

no you have to find x when y= 8

OpenStudy (anonymous):

0ok